OpenStudy (anonymous):

WILL MEDAL. Two marbles are to be pulled from a bag that contains 2 yellow marbles, 4 green marbles, and 6 orange marbles. After the first marble is drawn, it is not replaced What is the probability that both marbles will be the same color?

OpenStudy (mathmate):

Bag: 2Y 4G 6O, draw two without replacement. How many in all?

OpenStudy (anonymous):

12 in all @mathmate

OpenStudy (mathmate):

Good! Have you learned about "combinations", such as \(C^4_2\)?

OpenStudy (anonymous):

no

OpenStudy (mathmate):

Can you name me the outcomes that could be considered "success"?

OpenStudy (anonymous):

im not quite sure how to do this.

OpenStudy (mathmate):

Is YY a success, or is YG a success?

OpenStudy (anonymous):

yy?

OpenStudy (mathmate):

Success is considered the events that are included in our calculation of probabilities.

OpenStudy (mathmate):

Correct! Can you name all events that are successes?

OpenStudy (anonymous):

gg and oo?

OpenStudy (mathmate):

Perfect. Now you will have to calculate the probability of the three cases which lead to success, namely YY, GG and OO. So far so good?

OpenStudy (anonymous):

so would it be y * y and g*g or?

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

@sar12389 Say, what is the probability of getting a Y on the first draw?

OpenStudy (anonymous):

2/12

OpenStudy (anonymous):

4/12 g and 6/12 o @mathmate

OpenStudy (mathmate):

Good, have you done the multiplication law?

OpenStudy (anonymous):

im not sure

OpenStudy (mathmate):

Good, all three are correct for the _first_ draw. Now what is the probability of getting another Y after the first Y?

OpenStudy (anonymous):

1/12?

OpenStudy (anonymous):

3/12 g and 5/12 o

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

For YY, first one is 2/12, second is 1/11 (they do not replace). So by the multiplication law, the probability of YY is (2/12)*(1/11)=2/132. You can try the other two. The answer is the sum of all three, P(YY)+P(GG)+P(OO).

OpenStudy (anonymous):

so 43/132 ? what would the percent be @mathmate

OpenStudy (anonymous):

i have to choose from 79.5%, 81.5% , 85% or 90%

OpenStudy (anonymous):

i have to choose from 79.5%, 81.5% , 85% or 90%

OpenStudy (mathmate):

You would do a division of 43/132. For probability, I prefer to keep the fractions. Note: I think I got a different number. Can you check?

OpenStudy (anonymous):

i got 0.3257

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

That corresponds to 43/132. As I said, 0.3257 is a (rounded) approximately number. 43/132 is exact, which is preferred in probabilities. But I get a slightly different number than 43/132.

OpenStudy (anonymous):

what is the percent? @mathmate

OpenStudy (mathmate):

I get (2*1+4*3+6*5)/(12*11)=1/3=33% for two marbles of the same colour, and (2*10+4*8+6*6)/(12*11)=2/3=67% for two marbles of different colours. I do not know from where the choices came. All I can suggest is to reread the question to make sure there are no typos.